OpenStudy (anonymous):

Jason has two bags with 6 tiles each. The tiles in each bag are shown below: Make 6 squares. The squares are numbered sequentially from 1 to 6. Without looking, Jason draws a tile from the first bag and then a tile from the second bag. What is the probability of Jason drawing the tile numbered 5 from the first bag and an odd tile from the second bag? 3 over 6 4 over 6 3 over 36 4 over 36

OpenStudy (anonymous):

isnt this ques related to the chapter - probability

OpenStudy (anonymous):

no its math

OpenStudy (anonymous):

@pooja195 @bibby @EclipsedStar @AriPotta @freckles @Data_LG2 @camerondoherty @YanaSidlinskiy @Somy @Mashy @inowalst @Love_Ranaa @HelpBlahBlahBlah @linn99123 @matt101 @leonardo0430 @Jaynator495 @DullJackel09 @Mchilds15 @MakaylaTracy @andrewhaze @Teddyiswatshecallsme @Ilovecake @xo.minnie.xox @danish071996 @jgirl128 @ninaesb @Bookworm14 @aayushi.somani @nevermind_justschool @Elise_a18 @chycora @shortycme @iamabarbiegirl @IrishBoy123 @HaileyD @Brandonsnider188 @torifamelol @ameenbriscoe13 @skate-4-life270 @CallMeKiki @nobrainiachere @~SierraBozarth~ @tylermcmullen23 @nermaljean99 @jeffyblood @LXelle @youngnephew34 @Write0The0Right0For0Wrong0Doing @ashhhrodriguez @horsegirl325 @cvliboi @Dustinself42 @RasheedahOwl @Trvllest_Qveen110 @freegirl112 @deadpoolroxz @Spyro517 @StudyGuidance @NormaValenzuela

OpenStudy (anonymous):

@iGreen

OpenStudy (inowalst):

Please dont mass tag.

OpenStudy (anonymous):

sorry i really need help

OpenStudy (nobrainiachere):

It says this on my profile, please do not tag me in math because i can not help you, sorry

OpenStudy (anonymous):

anyone know anything?

OpenStudy (anonymous):

can someone help me tho?

OpenStudy (anonymous):

Hailey?

OpenStudy (anonymous):

@pooja195 @bibby @EclipsedStar @AriPotta @freckles @Data_LG2 @camerondoherty @YanaSidlinskiy @Somy @Mashy @Love_Ranaa @HelpBlahBlahBlah @linn99123 @matt101 @leonardo0430 @Jaynator495 @DullJackel09 @Mchilds15 @MakaylaTracy @andrewhaze @Teddyiswatshecallsme @Ilovecake @xo.minnie.xox @danish071996 @jgirl128 @ninaesb @Bookworm14 @aayushi.somani @nevermind_justschool @Elise_a18 @chycora @shortycme @iamabarbiegirl @IrishBoy123 @HaileyD @Brandonsnider188 @torifamelol @ameenbriscoe13 @skate-4-life270 @CallMeKiki @~SierraBozarth~ @tylermcmullen23 @nermaljean99 @jeffyblood @LXelle @youngnephew34 @Write0The0Right0For0Wrong0Doing @ashhhrodriguez @horsegirl325 @cvliboi @Dustinself42 @RasheedahOwl @Trvllest_Qveen110 @freegirl112 @deadpoolroxz @Spyro517 @StudyGuidance @NormaValenzuela

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

I have no clue about math

OpenStudy (anonymous):

why is @nobrainiachere here i didnt even tag him

OpenStudy (youngnephew34):

I'm not good at it

OpenStudy (nobrainiachere):

Two reasons, one you just tagged me -_- and two, i got a notification saying that you replied to this question.

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no its wrong wait let me resolve

OpenStudy (callmekiki):

Please don't tag me in math.

OpenStudy (anonymous):

"The tiles in each bag are shown below" i don't see anything .-. post a screenshot maybe?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@Deadpool117

OpenStudy (anonymous):

@pooja195 @bibby @EclipsedStar @AriPotta @freckles @Data_LG2 @camerondoherty @YanaSidlinskiy @Somy @Mashy @Love_Ranaa @HelpBlahBlahBlah @linn99123 @matt101 @leonardo0430 @Jaynator495 @DullJackel09 @Mchilds15 @MakaylaTracy @andrewhaze @Teddyiswatshecallsme @Ilovecake @xo.minnie.xox @danish071996 @jgirl128 @ninaesb @Bookworm14 @aayushi.somani @nevermind_justschool @Elise_a18 @chycora @shortycme @iamabarbiegirl @IrishBoy123 @HaileyD @Brandonsnider188 @torifamelol @ameenbriscoe13 @skate-4-life270 @CallMeKiki @~SierraBozarth~ @tylermcmullen23 @nermaljean99 @jeffyblood @LXelle @Write0The0Right0For0Wrong0Doing @ashhhrodriguez @horsegirl325 @cvliboi @RasheedahOwl @Trvllest_Qveen110 @freegirl112 @deadpoolroxz @Spyro517 @StudyGuidance @NormaValenzuela @uri @iGreen

OpenStudy (igreen):

No mass tagging.

OpenStudy (anonymous):

ERASE ME FROM THIS!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

STOP TAGGING ME OmL

OpenStudy (anonymous):

lol sorry

OpenStudy (spyro517):

count me in that as well

OpenStudy (igreen):

How many total tiles are there? @pandaluvs

OpenStudy (anonymous):

six

OpenStudy (anonymous):

well, 12 bc there are 2 bags

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then 12

OpenStudy (anonymous):

jeez ur gonna jam up the system by tagging so many people

OpenStudy (anonymous):

Sorry to sound rude

OpenStudy (anonymous):

well, no one ANSWERS!

OpenStudy (anonymous):

goodness.

OpenStudy (igreen):

Yes, there are 6 in one bag.

OpenStudy (yanasidlinskiy):

You do realize mass tagging is against the Code-of-Conduct.

OpenStudy (igreen):

How many 5's are in one bag?

OpenStudy (igreen):

@YanaSidlinskiy I already pointed that out.

OpenStudy (anonymous):

ok @iGreen and?

OpenStudy (igreen):

\(\color{blue}{\text{Originally Posted by}}\) @iGreen How many 5's are in one bag? \(\color{blue}{\text{End of Quote}}\) @pandaluvs

OpenStudy (yanasidlinskiy):

iGreen, I guess I overlooked. Thanks:) I just scrolled to the bottom of this post. Didn't bother reading that comments.

OpenStudy (anonymous):

one? @iGreen

OpenStudy (igreen):

Yep, so the probability for picking a 5 in one bag is 1/5.

OpenStudy (anonymous):

ok

OpenStudy (igreen):

How many odd tiles are in each bag?

OpenStudy (igreen):

@YanaSidlinskiy Okay.

OpenStudy (anonymous):

3

OpenStudy (igreen):

Sorry, made a mistake..that should be 1/6 for picking a 5.

OpenStudy (anonymous):

ok @iGreen

OpenStudy (igreen):

Yes, so the probability of picking an odd number from the 2nd bag is 3/6. Now multiply them together. 1/6 * 3/6 = ?

OpenStudy (anonymous):

0.08

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

As a fraction..

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

First off, 3/6 can be simplified to 1/2. \(\dfrac{1}{6} \times \dfrac{1}{2} \rightarrow \dfrac{1 \times 1}{6 \times 2}\)

OpenStudy (anonymous):

ITS 2/3

OpenStudy (anonymous):

@i

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

SOMEONE TELL ME IF IM RIGHT

OpenStudy (anonymous):

or ill tag you all again..... one .... by one.....

OpenStudy (anonymous):

@inowalst

OpenStudy (anonymous):

@ispike @Ilovecake @

OpenStudy (anonymous):

@shrutipande9

OpenStudy (deadpoolroxz):

wazagoinon?! i was sleeping ! X(

OpenStudy (anonymous):

lol

TheSmartOne (thesmartone):

What is \(\sf 6 \times 2+?\)

OpenStudy (igreen):

No.. it isn't..

TheSmartOne (thesmartone):

\(6\times 2=?\)**

OpenStudy (igreen):

1 * 1 = ?